40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 23:53:16 2019, gautam, Update, ALS, IR ALS locking attempt ALS_X_outOfLoopnoise.pdfALS_X_inLoopnoise.pdfCARM_loopShape.pdf
    Reply  Wed Jul 24 00:22:24 2019, gautam, Update, ALS, PSL frequency locked to XARM length using ALS ALS_X_noise_POX.pdfBR_comparison.pdfALS_CARM_OLG.pdf
       Reply  Thu Jul 25 12:25:56 2019, gautam, Update, ALS, IR ALSX noise ALS_broadPeaks.pdf
Message ID: 14800     Entry time: Mon Jul 22 23:53:16 2019     Reply to this: 14802
Author: gautam 
Type: Update 
Category: ALS 
Subject: IR ALS locking attempt 

Summary:

My goal tonight was to lock the PSL frequency to be resonant in the XARM cavity, using the PSL+EX beat as the error signal. I was not successful - mainly, I was plagued by huge BR mode coupling in the error signal, and I could not enable the BR notch filter in the control loop without breaking the lock. Need to think about next steps.

Details:

  • POX and POY locking was easily restored.
  • EX green alignment was tweaked at the end-table. A large YAW correction was required, which I opted to apply on the mechanical mirror mounts rather than the PZTs. GTRX ~0.4 was recovered.
  • The arm cavity length was first locked using POX as an error signal 
    • Then I looked at the out-of-loop ALS noise, trusting the DFD's V/Hz calibration (red-trace in Attachment #1).
    • I judged it to be close enough to the benchmark reference (green-trace in Attachment #1), and so decided that I could go ahead and try locking.
  • A modified version of the script /opt/rtcds/caltech/c1/scripts/XARM/Lock_ALS_XARM.py was used to transition control from POX to the ALS error signal
    • I found that I had to change the sign of the CARM loop gain for the servo to remain stable (in this config, CARM-->MC2 length, thereby modifying the IMC frequency to keep the PSL resonant in the XARM cavity).
    • I don't know why this sign change was required - we are still sticking to the same convention that the beat frequency increases when the temperature slider for the EX laser is incremented in counts.
    • The script failed multiple times at the BOOST/BR notch filter enabling step.
    • Doing these steps manually, I found that turning the BR notch, FM6, ON destroyed the lock immediately.
    • Motivated by this observation, I looked at the in-loop error signal spectrum, see Attachment #2. Here, the PSL frequency is servoed by the ALS error signal, but the BR notch filter isn't enabled.
    • The Bounce-mode peak is huge - where is this coming from? It is absent in the ALS spectrum when the XARM is locked with POX. So it is somehow connected with actuating on the MC2 suspension? Or is it that the FM6=BounceRoll filter of the XARM loop is squishing the noise when looking at the ALS spectrum in POX lock, i.e. Attachment #1? In which case, why can't I engage FM6 for the CARM loop???

Anyway, now that I have a workable set of settings that gets me close to the ALS lock of the XARM, I expect debugging to proceed faster.


Update 2019 July 23: I looked at the control loop shape today, see Attachment #3. I'm not sure I understand why the "BounceRoll" filter  in this filter bank looks like a resonant gain rather than a notch, as it does for the Oplev or SUSPOS loops for example - don't we want to not actuate at these frequencies because the reason the signal exists is because of the imperfect OSEM/magnet positioning? This does not explain the spectrum shown in Attachment #2 however, as that filter was disabled.

Attachment 1: ALS_X_outOfLoopnoise.pdf  40 kB  Uploaded Tue Jul 23 01:12:44 2019  | Hide | Hide all
ALS_X_outOfLoopnoise.pdf
Attachment 2: ALS_X_inLoopnoise.pdf  41 kB  Uploaded Tue Jul 23 01:12:52 2019  | Hide | Hide all
ALS_X_inLoopnoise.pdf
Attachment 3: CARM_loopShape.pdf  17 kB  Uploaded Tue Jul 23 16:11:38 2019  | Hide | Hide all
CARM_loopShape.pdf
ELOG V3.1.3-